The demand for ice cream cones is p=1600 and Qd is 2. The supply of ice cream cones is p= 400 and Qs if 1.The price of a cone is expressed in cents ,and the quaquantities are expressed in cones per day.To find the equilibrium price (P*) and the equilibrium quantity (Q*) substitute Q* (for QD and QS and p.for QD, P=a-bQd and for Qs P=c+dQS)
1
Expert's answer
2019-04-01T11:06:38-0400

Qd :

P=a-bQd

bQd=a-P

1600 cents=16 dollars

2=a-16

a=18

Qd=18-P


Qs:

P=c+d*Qs

dQs=P-c

400 cents=4 dollars

1=4-c

c=3

Qs=P-3


At equilibrium, Qd = Qs

18-P=P-3

2P=21

P*=10,5 $

Q*=18-10,5=7,5

equilibrium price is 10,5 $ or 1050 cents

equilibrium quantity is 7,5 items